# Università della Svizzera italiana, Switzerland

### v1

@@ Abstract
Large Language Models (LLMs) have shown an impressive capability in code generation and, specifically, to automatically implement requirements described in natural language. The LLM effectiveness generally increases with its size: The higher the number of LLM’s trainable parameters the better its ability to implement code. However, when it comes to deploying LLM-based code generators, larger LLMs pose significant challenges related to their memory (and, consequently, carbon) footprint. A previous work by Wei et al.proposed to leverage quantization techniques to reduce the memory footprint of LLM-based code generators without substantially degrading their effectiveness. In short, they studied LLMs featuring up to 16B parameters, quantizing their precision from floating point 32 bits down to int 8 bits and showing their limited impact on code generation performance. Given the fast pace at which LLM capabilities and quantization techniques are evolving, in this work we present a differentiated replication of the work by Wei et al.in which we consider (i) on the one side, more recent and larger code-related LLMs, of up to 34B parameters; (ii) the latest advancements in model quantization techniques, which allow pushing the compression to the extreme quantization level of 2 bits per model parameter and; (iii) different types of calibration datasets to guide the quantization process, including code-specific ones. Our empirical evaluation reveals that the new frontier for LLM quantization is 4-bit precision, resulting in an average memory footprint reduction of 70% compared to the original model without observing any significant decrease in performance. Additionally, when the quantization becomes even more extreme (3 and 2 bits), a code-specific calibration dataset helps to limit the loss of performance.
Keywords:
Empirical Software Engineering Pre-Trained Models AI For Software Engineering Green AI1 Introduction
The advent of deep learning (DL) in software engineering pushed the boundaries of automated code generation, with techniques and tools able to implement complete software pieces of functionality starting from a natural language description. The most recent and advanced code generators are grounded on Large Language Models (LLMs), namely DL models featuring billion parameters. Besides closed commercial software-specific tools such as GitHub CoPilot (cop, 2021) and general purpose ones such as GPT4 (Achiam et al., 2023), there also exist several open-source LLMs specifically pre-trained on code-related data. These include CodeLlama (Roziere et al., 2023) and DeepSeek Coder (Guo et al., 2024), both being families of code-specialized LLMs of different sizes.
@@
While LLMs opened up new frontiers in code generation, their usage has downsides. On the one hand, developers could leverage closed models, often available through APIs and often also integrated into the development environments. This is the case, for example, of GitHub CoPilot (cop, 2021) or GPT4 (Achiam et al., 2023). While in such cases, the training and inference are performed on third-party servers, using such models introduces data privacy issues (Novelli et al., 2024; Wu et al., 2024; Huang et al., 2023). Also, such models are general-purpose and not customized to work on specific, proprietary technology. An alternative would be to leverage open-source models by training and using them locally. However, the large number of trainable parameters implies a conspicuous cost and environmental impact during the training phase. Also, such models impose hardware requirements for the inference phase (i.e., when the LLM is used to generate recommendations). For example, DeepSeek Coder-33B would require, at inference time, 64GB of memory (VRAM), which is more than what most of modern laptops are equipped with. Certainly, this solution could be circumvented by deploying models for inference on local servers and using them as services, yet this still implies costs, increased complexity, and network overhead.
@@
Wei et al. (2023a) have been the first investigating code generation via quantization, showing that by quantizing the weights of a 16B-parameters LLM (i.e., CodeGen (Nijkamp et al., 2023)) to int8 it was possible to reduce its memory footprint by 29% with a negligible reduction in the code generation capabilities. Due to the fast pace at which LLM capabilities and quantization techniques are evolving, in this work we present a differentiated replication of the study conducted by Wei et al. (2023a). In particular, we experiment with the latest advancements in model quantization techniques, which allow us to push the compression to extreme quantization levels such as 2 bits per model parameter. Also, we consider different types of calibration datasets to guide the quantization process (i.e., the training data used to approximate the original weights), including code-specific ones. Last but not least, we focus on newer and larger LLM-based code generators, up to 34B parameters, and extend the study to two programming languages (Python and Java) instead of Python-only (Wei et al., 2023a).
@@
We consider two families of code-specific LLMs, namely CodeLlama (Roziere et al., 2023) and DeepSeek Coder (Guo et al., 2024), experimenting with three of their variants: CodeLlama 7B, 13B, and 33B, and DeepSeek Coder 1.3B, 7B, and 33B. We apply on them the recently proposed Additive Quantization with Learned Multi-Codebooks (AQLM) (Egiazarian et al., 2024) quantization techniques, which achieved new state-of-the-art results for NLP benchmarks at extreme quantization levels (e.g., 2 bits). In particular, we experiment with quantizations at 8, 4, 3, and 2 bits. We assess the code generation capabilities of the baseline (i.e., the original model) and of the quantized models on two benchmarks, MultiPL-E (Cassano et al., 2023) and McEval (Chai et al., 2024). We focus on the Java and Python instances from the two benchmarks for a total of 414 implementation problems. We report the percentage of cases in which the subject models managed to correctly implement the required code at the first attempt. On top of that, we experiment with three different calibration datasets: (i) a general-purpose dataset with randomly sampled content (referred to as the random dataset); (ii) a mixed dataset comprising 50% GitHub code files and 50% samples from Stack Overflow discussions; and (iii) a code-specific dataset containing only Java and Python code.
@@
The obtained results show that, thanks to novel quantization techniques, such as AQLM, it is possible to quantize LLMs for the task of code generation down to 4 bits without observing any significant decrease in performance. Such a process results in a memory footprint reduced by 70%, allowing the deployment of very large models also on hardware-constrained devices (and, in general, lowering the deployment cost). Interestingly, the impact of extreme quantization levels on software-related practices has yet to be explored. To the best of our knowledge, we are the first to demonstrate the feasibility of reducing precision below 4 bits, going as low as 3 and 2 bits, across multiple large code models and benchmarks. Yet, this results in a decrease in performance, though it can be mitigated by adopting a calibration dataset that includes code. Finally, we found that larger code models are more resilient to information loss during extreme quantization (e.g., 2-bit quantization).
@@
Our study, on the one side, confirms the findings by Wei et al. (2023a) with larger models, multiple programming languages, and different quantization techniques. On the other side, and more importantly, it shows that the recent progress makes quantization even more attractive, due to the potential for extreme quantization. By pushing the precision down to 3 and 2 bits, we are laying the groundwork for more efficient code-specialized LLMs. This, in turn, reduces computational costs and memory footprint, with a trade-off in terms of performance that remains manageable.
We release all code and data used in our study, as well as the quantized models (Giagnorio et al., 2025).
The paper is organized as follows. Section 2 provides background notions and discusses related literature concerning on the one hand the use of LLMs in software engineering and, on the other hand, techniques to improve the efficiency of deep learning models. Section 3 defines the study and details its planning and analysis methodology. Results are reported and discussed in Section 4. Then, Section 5 outlines implications of our finding for both research and practice, while the study’s threats to validity are discussed in Section 6. Finally, Section 7 concludes the paper and outlines directions for future work.
2 Background and Related Work
In this section, we start (Section 2.1) by highlighting the multifaceted capabilities of LLMs in handling various complex tasks, particularly those related to software engineering, such as bug-fixing and code generation. Next (Section 2.2), we review the relevant literature concerning, on the one hand, the environmental impact of machine learning models and, on the other hand, techniques to make models more efficient—especially in the context of software engineering—and consequently more environment-friendly. In particular, we focus on model compression techniques. Finally (Section 2.3), we explain the quantization technique employed in our study, specifically AQLM.
2.1 Large Language Models for Software Engineering
LLMs are a specialized subset of neural network models engineered to understand, manipulate, and generate natural language. The remarkable capabilities of these AI-driven models arise from their extensive pre-training on vast amounts of textual data from diverse sources, combined with the large number of parameters they encompass. Notable examples in this field include GPT-4 (Achiam et al., 2023), Claude (Anthropic, 2024), and Gemini (Team et al., 2023), all of which boast over 100 billion parameters.
LLMs have been successfully applied across various areas in software engineering including, but not limited to, requirements analysis (Marques et al., 2024), code generation (Li et al., 2022; Mastropaolo et al., 2024), code translation (Fan et al., 2023), code summarization (Haldar and Hockenmaier, 2024; Ahmed et al., 2024; Ahmed and Devanbu, 2022; Sun et al., 2024a), and test case generation (Tufano et al., 2020). While general-purpose LLMs like GPT-4 and Claude can support various domains, researchers have also developed domain-specific models tailored to tackle software engineering activities and code-related tasks. For instance, open-source LLMs such as InCoder (Fried et al., 2023), CodeLlama (Roziere et al., 2023), StarCoder (Li et al., 2023b), DeepSeek-Coder (Guo et al., 2024) and SantaCoder (Allal et al., 2023) have been designed specifically to handle these tasks, offering specialized capabilities that can substantially automate software-related tasks.
2.2 Efficiency and Environmental Impact of AI models for SE
@@
The deployment of billion-parameter LLMs imposes substantial computational demands and considerable energy requirements, often reliant on non-renewable resources, sparking concerns about environmental sustainability. Recently, Castaño et al. (2023) documented the correlation between the DL models’ carbon footprint and various factors like model size and parameter count. The significant energy consumption and carbon emissions associated with LLMs have also been widely documented (Patterson et al., 2021; Strubell et al., 2020), emphasizing the urgent need for techniques aimed at making these models energy-efficient. In the remainder of this section, we highlight recent advancements in Green AI and how these are being translated into the software engineering domain.
Parameter-Efficient Fine-Tuning (PEFT) optimizes fine-tuning by updating only a subset of the model’s parameters. This can be done by introducing additional layers (Houlsby et al., 2019), decomposing weight gradients into specialized matrices (Hu et al., 2022), or via prompt-tuning techniques (Lester et al., 2021; Li and Liang, 2021). PEFT has been applied in tasks like code generation and code summarization. For example, Wang et al. (2023) introduced Adapter tuning for code search and summarization, while Ayupov and Chirkova (2022) evaluated the effectiveness of two PEFT methods, Adapters (Wang et al., 2023) and LoRA (Hu et al., 2022), for tasks such as code generation and summarization, reporting significant performance gains. Weyssow et al. (2023) explored four PEFT techniques: LoRA, IA3 (Liu et al., 2022), prompt tuning (Lester et al., 2021), and prefix tuning (Li and Liang, 2021) — in the context of code generation. Another study (Liu et al., 2023) compared Adapter, LoRA, prefix tuning, and MHM (He et al., 2022) PEFT methods for code generation and summarization. The above-discussed works report positive findings, showing the effectiveness of PEFT even when compared to a classic fine-tuning updating all model parameters.
Knowledge Distillation (KD) belongs to the class of model compression techniques and aims at teaching a smaller model to emulate a larger one (Hsieh et al., 2023). Noteworthy KD applications in code-related models include CodeAlpaca (Chaudhary, 2023), Magicoder (Wei et al., 2023b), WaveCoder (Yu et al., 2023), and MFTCoder (Liu et al., 2023), most of which focus on code generation. Studies like the one by Su and McMillan (2024) have shown that distilled models can match the performance of much larger counterparts like GPT-3.5 in specific tasks (e.g., code summarization) without significant losses in performance metrics. Although KD unlocks the possibility of using smaller models mimicking the capabilities of a larger one, an expensive training phase remains necessary.
Quantization is a model compression technique that aims at reducing the model’s memory footprint by representing its parameters (weights and/or activations) in a lower-precision format, such as 8-bit integers, instead of the standard 16-bit or 32-bit floating points (Gholami et al., 2022). Quantization can be applied through two possible strategies: Quantization-Aware Training (QAT) and Post-Training Quantization (PTQ). QAT (Esser et al., 2020) integrates quantization during training, necessitating a complete model retraining. Specific adaptations for LLMs, such as LLM-QAT (Liu et al., 2024) and EdgeQAT (Shen et al., 2024), can usually be applied to LLMs having limited size. Indeed, the high GPU time and memory demands render QAT impractical for large-scale LLMs, making PTQ a preferable alternative. PTQ (Cai et al., 2020) converts pre-trained models to fixed-point networks without revisiting the initial training stages. This method requires the model parameters to be adjusted to minimize quantization errors using a calibration dataset, but such an “additional training” is quite cheap to run.
PTQ techniques can be further divided into weights-only and weight-activation quantization. Weights-only quantization, such as GPTQ (Frantar et al., 2022) and strategies in PB-LLM (Yuan et al., 2024) and SpQR (Dettmers et al., 2024), focuses on minimizing precision loss by adjusting weight bit-widths and applying scale transformations for critical weight channels. Weight-activation quantization, exemplified by SmoothQuant (Xiao et al., 2023) and LLM.int8() (Dettmers et al., 2022), compresses both weights and activations, utilizing techniques like mixed-precision decomposition and channel-wise scaling. For ultra-low bit quantization (), advanced PTQ methods such as QuIP# (Tseng et al., 2024) and Additive Quantization of Language Models (AQLM) (Egiazarian et al., 2024) have been developed, with the latter outperforming QuIP# and achieving Pareto-optimal performance (i.e., maximization of accuracy for a given model’s size) in extreme compression scenarios (less than 3 bits per parameter). This is done thanks to the application of a novel technique named MCQ (Multi-Codebook Quantization). These techniques have been tested on LlaMA (Touvron et al., 2023) and OPT (Zhang et al., 2022) on natural language-related benchmarks.
In the software engineering domain, little research has been conducted in this direction.
Sun et al. (2024b) studied Dynamic Inference to optimize the code completion task and improve LLMs efficiency. Their proposed technique (Stop&Exit Controller) skips, on average, 1.7 layers of the analyzed LLMs, improving the inference time by 11.2% with negligible accuracy loss. Differently from this study, we target the code generation task and aim at reducing the model’s memory footprint while preserving the original performance.
d’Aloisio et al. (2024) investigated the effects of several compression techniques (knowledge distillation, quantization, and pruning) on code tasks such as vulnerability detection, code summarization, and code search. Their findings show that, while these techniques can help to reduce computational costs and inference time, they may result in significant performance loss. Quantization, instead, is the most promising technique for achieving comparable accuracy to the original model at lower memory costs. Our study explores more in-depth the quantization technique on the code generation task and employs the state-of-the-art technique for lowering the model precision down to 2 bits.
The most related study is the investigation by Wei et al. (2023a), which inspired our work. In this research, the authors explored 8-bit quantization of code models (PLBART, Code-T5, InCoder, CodeGen) for code generation. Their findings revealed that 8-bit quantized versions of the CodeGen and InCoder models achieved improved energy efficiency accompanied by a limited drop in performance, particularly for the largest model featuring 16B parameters. Our study looks at new state-of-the-art code models - CodeLlama and DeepSeek-Coder (up to 34B parameter models), and employs AQLM for extreme quantization down to 2 bits. Further details about AQLM are discussed in Section 2.3. Moreover, as recent research (Williams and Aletras, 2024) highlights the critical role of calibration data in determining model performance, our study also investigates the effect of different calibration datasets on the quantization outcomes of these large code models. As documented later, we experiment with a variety of calibration datasets, including those that combine code and natural language, showing their benefits at extreme quantization levels (3 and 2 bits).
2.3 Additive Quantization of Language Models (AQLM)
@@
For compressing LLM weights, the “direct” quantization method involves selecting an appropriate quantization grid and normalization for each subcomponent of the matrix. Subsequently, weights are mapped onto this grid either through direct rounding (Round-To-Nearest) (Dettmers et al., 2022) or through more sophisticated methods. In contrast, AQLM employs Multi-Codebook Quantization (MCQ): A generalization of vector quantization (VQ) used in the approximate nearest neighbor search algorithms quantizes multiple vector dimensions simultaneously by learning “codebooks”—sets of learnable candidate vectors for data encoding. In VQ, a database vector is divided into sub-groups, with each group encoded using a vector from a learned codebook. This process facilitates efficient computation of distances or dot products for similarity searches by exploiting the linearity of dot products. Additive Quantization (AQ) (Babenko and Lempitsky, 2014), a popular MCQ algorithm, performs vector optimization in a way that for a linear layer with input and output features, characterized by its weight matrix and a set of calibration inputs , the objective is to find a configuration of quantized weights that minimizes the squared error between the outputs of the original and compressed layers. To do so, AQ divides weight rows into groups of consecutive elements and represents each group by summing vectors selected from multiple codebooks , where each codebook contains vectors for -bit encoding. Each group’s weight is quantized by combining a chosen vector from each codebook, encoded as a one-hot vector .
The full-weight representation is then achieved by concatenating these quantized segments according to the following structure:
@@
where denotes concatenation and each is a one-hot code that specifies the selection from the -th codebook for the -th output unit and the -th group of input dimensions. With these techniques, AQLM ensures the integrity of the output from each layer and Transformer block.
3 Design

### v2

# Evaluating the Impact of Post-Training Quantization on Large Language Models for Code Generation

@@ Abstract.
Large Language Models (LLMs) have shown an impressive capability in code generation. The LLM effectiveness generally increases with its size: The higher the number of LLM’s trainable parameters the better its ability to implement code. However, when it comes to deploying LLM-based code generators, larger LLMs pose significant challenges related to their memory (and, consequently, carbon) footprint. A previous work by Wei et al. proposed to leverage quantization techniques to reduce the memory footprint of LLM-based code generators without substantially degrading their effectiveness. In short, they studied LLMs featuring up to 16B parameters, quantizing their precision from floating point 32 bits down to int 8 bits and showing their limited impact on code generation performance. Given the fast pace at which LLM capabilities and quantization techniques are evolving, in this work we present a differentiated replication of the work by Wei et al. in which we consider (i) on the one side, more recent and larger code-related LLMs, of up to 34B parameters; (ii) the latest advancements in model quantization techniques, which allow pushing the compression to the extreme quantization level of 2 bits per model parameter and; (iii) different types of calibration datasets to guide the quantization process, including code-specific ones. Our empirical evaluation reveals that the new frontier for LLM quantization is 4-bit precision, resulting in an average memory footprint reduction of 70% compared to the original model without observing any significant decrease in performance. Additionally, when the quantization becomes even more extreme (3 and 2 bits), a code-specific calibration dataset helps to limit the loss of performance.
1. Introduction
@@
Large Language Models (LLMs) opened up new frontiers in code generation. However, their usage also introduced new concerns. Developers could leverage closed models, often available through APIs and often also integrated into the development environments. This is the case, for example, of GitHub CoPilot (23) or GPT (Achiam et al., 2023). While in such cases, the training and inference are performed on third-party servers, using such models introduces data privacy issues (Novelli et al., 2024; Wu et al., 2024; Huang et al., 2023). Also, such models are general-purpose and not customized to work on specific, proprietary technology. An alternative would be to leverage open models by training and using them locally. However, the large number of trainable parameters implies a conspicuous cost and environmental impact during the training phase. Also, such models impose hardware requirements for the inference phase (i.e., when the LLM is used to generate recommendations). For example, DeepSeek Coder-33B would require, at inference time, 64GB of memory (VRAM), which is more than what most of modern laptops are equipped with. Certainly, this solution could be circumvented by deploying models for inference on local servers and using them as services, yet this implies costs, increased complexity, and network overhead.
@@
Wei et al. (Wei et al., 2023) have been the first investigating code generation via quantization, showing that by quantizing the weights of a 16B-parameters LLM (i.e., CodeGen (Nijkamp et al., 2023)) to int8 it was possible to reduce its memory footprint by 29% with a negligible reduction in the code generation capabilities. Due to the fast pace at which LLM capabilities and quantization techniques are evolving, in this work we present a differentiated replication of the study conducted by Wei et al. (Wei et al., 2023). We experiment with the latest advancements in model quantization techniques, which allows to push the compression to extreme quantization levels such as 2 bits per parameter. Also, we consider different types of calibration datasets to guide the quantization process (i.e., the training data used to approximate the original weights), including code-specific ones. Last but not least, we focus on newer and larger LLM-based code generators, up to 34B parameters, and extend the study to two programming languages (Python and Java) instead of Python-only (Wei et al., 2023).
@@
We consider two families of code-specific LLMs, namely CodeLlama (Roziere et al., 2023) and DeepSeek Coder (Guo et al., 2024), experimenting with three of their variants: CodeLlama 7B, 13B, and 33B, and DeepSeek Coder 1.3B, 7B, and 33B. We apply on them the recently proposed Additive Quantization with Learned Multi-Codebooks (AQLM) (Egiazarian et al., 2024) quantization techniques, which achieved new state-of-the-art results for Natural Language Processing (NLP) benchmarks at extreme quantization levels (e.g., 2 bits). In particular, we experiment with quantizations at 8, 4, 3, and 2 bits. We assess the code generation capabilities of the baseline (i.e., the original model) and of the quantized models on two benchmarks, MultiPL-E (Cassano et al., 2023) and McEval (Chai et al., ). We focus on the Java and Python instances from the two benchmarks for a total of 414 implementation problems. We report the percentage of cases in which the subject models managed to correctly implement the required code at the first attempt. On top of that, we experiment with three different calibration datasets: (i) a general-purpose dataset with randomly sampled content (random dataset); (ii) a mixed dataset comprising 50% GitHub code files and 50% samples from Stack Overflow discussions; and (iii) a code-specific dataset containing only Java and Python code.
@@
The obtained results show that, thanks to novel quantization techniques, such as AQLM, it is possible to quantize LLMs for the task of code generation down to 4 bits without any significant decrease in performance. Such a process results in a memory footprint reduced by 70%, allowing the deployment of very large models also on hardware-constrained devices (and, in general, lowering the deployment cost). To the best of our knowledge, we are the first to demonstrate the feasibility of reducing precision below 4 bits, going as low as 3 and 2 bits, across multiple large code models and benchmarks. Indeed, for the first time, we show that using ad-hoc calibration datasets, especially those containing code samples and technical language, helps mitigating the performance degradation typically associated with such extreme quantization levels. Finally, we found that larger code models are more resilient to information loss during extreme quantization.
@@
Our study, on the one side, confirms the findings by Wei et al. (Wei et al., 2023) with larger models, multiple programming languages, and different quantization techniques. On the other side, and more importantly, it shows that the recent progress makes quantization even more attractive, due to the potential for extreme quantization. By pushing the precision down to 3 and 2 bits, we are laying the groundwork for more efficient code-specialized LLMs and the comprehension of the limits of quantization techniques in this context. All the code and data used in our study are publicly available (43).
2. Background and Related Work
Green AI for Software Engineering
@@ 2. Background and Related Work
Castaño et al. (Castaño et al., 2023) documented the correlation between the DL models’ carbon footprint and various factors like model size and parameter count. The significant energy consumption and carbon emissions associated with LLMs have also been widely documented (Patterson et al., 2021; Strubell et al., 2020), emphasizing the need for techniques aimed at making these models energy-efficient.
Several approaches have been proposed in this direction, such as Parameter-Efficient Fine-Tuning (PEFT) (Houlsby et al., 2019; Hu et al., 2022), Knowledge Distillation (KD) (Hsieh et al., 2023; Su and McMillan, 2024), and Quantization (Gholami et al., 2022).
PEFT and KD typically require an additional training phase, which remains very expensive for large models.
Quantization reduces the memory footprint of large language models (LLMs) by representing model parameters in lower-precision formats, such as 8-bit integers, instead of standard 16- or 32-bit floating points (Gholami et al., 2022). Two main strategies exist: Quantization-Aware Training (QAT) and Post-Training Quantization (PTQ). While QAT integrates quantization during model training (Esser et al., 2020) and has been adapted for smaller LLMs (Liu et al., 2024b), its high computational cost makes it impractical for large-scale models. PTQ, in contrast, converts pre-trained models to lower-precision formats with minimal training, typically using lightweight calibration datasets (Cai et al., 2020). Due to these advantages, we focus only on PTQ techniques in this work.
PTQ approaches include weights-only quantization (Frantar et al., 2022; Yuan et al., 2024; Dettmers et al., 2024), and weight-activation quantization (Xiao et al., 2023; Dettmers et al., 2022). For more aggressive compression () bit, recent methods like QuIP# (Tseng et al., 2024), VPTQ (Liu et al., 2024a) and AQLM (Egiazarian et al., 2024) have advanced the state of the art. AQLM, in particular, achieves Pareto-optimal performance in sub-3-bit scenarios through Multi-Codebook Quantization (MCQ), significantly outperforming prior techniques on natural language-related benchmarks.
In the software engineering domain, little research has been conducted in this direction. Recent studies have explored model compression techniques to enhance the efficiency of large language models (LLMs) for code-related tasks. Sun et al. (2024b) proposed a dynamic inference method (Stop&Exit Controller) that accelerates code completion by skipping layers with minimal performance degradation. QLoRA, which combines 4-bit quantization with LoRA fine-tuning, has shown promise for code generation and summarization tasks (Weyssow et al., 2023; Afrin et al., 2025). A broader evaluation of compression techniques—including pruning, distillation, and quantization—by d’Aloisio et al. (2024) confirmed quantization as the most effective for maintaining performance while reducing memory usage.
The most relevant prior work is by Wei et al. (2023), who investigated 8-bit quantization of large code models such as PLBART, CodeT5, and CodeGen, demonstrating improved energy efficiency with only marginal performance loss. Building on this, our study targets state-of-the-art models—CodeLlama and DeepSeek-Coder (up to 34B parameters)—and applies AQLM for extreme quantization down to 2 bits. Furthermore, motivated by findings from Williams and Aletras (2024) on the influence of calibration data, we systematically evaluate different calibration datasets to assess their role in preserving model quality under aggressive quantization.
Additive Quantization of Language Models
@@
Additive Quantization of Language Models (AQLM) employs Multi-Codebook Quantization (MCQ): A generalization of vector quantization (VQ) used in the approximate nearest neighbor search algorithms quantizes multiple vector dimensions simultaneously by learning “codebooks”—sets of learnable candidate vectors for data encoding. In VQ, a database vector is divided into sub-groups, with each group encoded using a vector from a learned codebook. This process facilitates efficient computation of distances or dot products for similarity searches by exploiting the linearity of dot products. Additive Quantization (AQ) (Babenko and Lempitsky, 2014), a popular MCQ algorithm, performs vector optimization in a way that for a linear layer with input and output features, characterized by its weight matrix and a set of calibration inputs , the objective is to find a configuration of quantized weights that minimizes the squared error between the outputs of the original and compressed layers. AQ divides weight rows into groups of consecutive elements and represents each group by summing vectors selected from multiple codebooks , where each codebook contains vectors for -bit encoding. Each group’s weight is quantized by combining a chosen vector from each codebook, encoded as a one-hot vector . The full-weight representation is then achieved by concatenating these quantized segments according to the following structure:
@@
where denotes concatenation and each is a one-hot code that specifies the selection from the -th codebook for the -th output unit and the -th group of input dimensions.
3. Design